OpenStudy (anonymous):

Calculus Find the length of the parabolic segment r=14/(1+cos(theta)), 0<=theta<=pi/2

OpenStudy (anonymous):

I know I need to use this formula, but I'm not sure how: \[L=\int\limits_{0}^{\pi/2}\sqrt{r^2 +(dr/d \Theta)^2}\]

zepdrix (zepdrix):

Hmmm the algebra isn't working out for me. Let's try to take this slow and see what's going on.

zepdrix (zepdrix):

So at some point we'll need our r'. Are you able to determine the derivative of your function r?

OpenStudy (anonymous):

Even that step is giving me trouble.

zepdrix (zepdrix):

Writing it like this might help, \[\Large\rm r=\frac{14}{1+\cos \theta}=14(1+\cos \theta)^{-1}\]Allows us to apply the power rule, into chain rule.

zepdrix (zepdrix):

Alternatively, you can apply the quotient rule, but that's not very fun when your numerator is just a constant.

zepdrix (zepdrix):

Still too hard? :3

zepdrix (zepdrix):

\[\Large\rm r'=-14(1+\cos \theta)^{-2}\color{royalblue}{\left(1+\cos \theta\right)'}\]We apply our power rule, then set up our chain rule. We need to multiply by the derivative of the inner function.

zepdrix (zepdrix):

\[\Large\rm r'=-14(1+\cos \theta)^{-2}\color{orangered}{\left(-\sin \theta\right)}\]

zepdrix (zepdrix):

\[\Large\rm \frac{dr}{d\theta}=\frac{14\sin \theta}{(1+\cos \theta)^2}\]Something like that, yes? :o

OpenStudy (anonymous):

ok. Just got there, working it out on paper

OpenStudy (anonymous):

Now I have \[L= \int\limits_{0}^{\pi/2}\sqrt{14/\cos(\theta) + (14\sin(\theta))/(1+\cos(\theta))^2}\]

zepdrix (zepdrix):

Woops, we want to square our r as well, yes?

zepdrix (zepdrix):

Lemme fix that up real quick:

zepdrix (zepdrix):

\[\Large\rm L= \int\limits\limits_{0}^{\pi/2}\sqrt{\left[\frac{14}{(1+\cos \theta)}\right]^2 + \left[\frac{14\sin \theta}{(1+\cos \theta)^2}\right]^2}~d\theta\]Mmmm like this I think?

OpenStudy (anonymous):

Nice! looks much cleaner and has the squares I left out

zepdrix (zepdrix):

The algebra gets a little nasty here.... hmm

zepdrix (zepdrix):

\[\Large\rm L= 14\int\limits_{0}^{\pi/2}\sqrt{\frac{(1+\cos \theta)^2}{(1+\cos \theta)^4} + \frac{\sin^2\theta}{(1+\cos \theta)^4}}~d\theta\]Lemme know if you're confused by this step. I did a few different things here all at once.

zepdrix (zepdrix):

I factored the 14 out of the root, I applied the squares, then I multiplied the first term by (1+cos theta)^2 on top and bottom to get a common denominator.

OpenStudy (anonymous):

Ok, I follow.

OpenStudy (anonymous):

Now the numerators can be on the same line and we can take the square root?

zepdrix (zepdrix):

Yes, good good. Combine fractions, take the square root of the bottom and pull it out,\[\Large\rm L= 14\int\limits\limits_{0}^{\pi/2}\frac{1}{(1+\cos)^2}\sqrt{(1+\cos \theta)^2+\sin^2\theta}~~d\theta\]

zepdrix (zepdrix):

Still have some work to do with the top though, keep in mind you can't take the root of the top,\[\Large\rm \sqrt{a^2+b^2}\ne a+b\]

OpenStudy (anonymous):

Is there a trig equivalency for this part? \[(1 + \cos \theta )^2\]

zepdrix (zepdrix):

No, we have to expand it out.

zepdrix (zepdrix):

We'll have a nice simplification after expanding it though.

zepdrix (zepdrix):

\[\large\rm L= 14\int\limits\limits\limits_{0}^{\pi/2}\frac{1}{(1+\cos)^2}\sqrt{(1+2\cos \theta+\cos^2\theta+\sin^2\theta}~~d\theta\]Somethingggggg like that, yah?

OpenStudy (anonymous):

okay. so under the square root turns to this?\[2+2\cos \theta\]

zepdrix (zepdrix):

Mmmm ok good. Factoring a 2 out of each term and pulling it out of the root gives us,\[\large\rm L= 14\sqrt{2}\int\limits_{0}^{\pi/2}\frac{1}{(1+\cos)^2}\sqrt{1+\cos \theta}~~d\theta\]

OpenStudy (anonymous):

Now a substitution for 1+cos?

zepdrix (zepdrix):

Hmmm :( Here's where I'm getting stuck. We can divide out sqrt{1+cos(theta)} from top and bottom, giving us,\[\large\rm L= 14\sqrt{2}\int\limits_{0}^{\pi/2}\frac{1}{(1+\cos)^{3/2}}~d\theta\]Buttttttt, we don't have a nice u-sub we can apply. See how there is no Sin(theta) showing up anywhere? We would have trouble with our du.

zepdrix (zepdrix):

I'm gonna throw it into wolfram a sec, to try and see where I goofed up.

zepdrix (zepdrix):

Hmm I can't figure it out ;c @aum @dan815 @jim_thompson5910 @ganeshie8

OpenStudy (anonymous):

Whats wrong with making u = (1 +cos theta)?

zepdrix (zepdrix):

\[\Large\rm du=-\sin \theta~d \theta\]That sine is a little difficult to deal with.

ganeshie8 (ganeshie8):

\(\large 1+\cos\theta = 2\cos^2(\theta/2)\)

ganeshie8 (ganeshie8):

\[\large\begin{align}\rm L&= 14\sqrt{2}\int\limits_{0}^{\pi/2}\frac{1}{(1+\cos)^{3/2}}~d\theta \\~\\&= 14\sqrt{2}\int\limits_{0}^{\pi/2}\frac{1}{(2\cos^2(\theta/2))^{3/2}}~d\theta\\~\\&=7 \int\limits_{0}^{\pi/2} \sec^{3}(\theta/2)~d\theta\\~\\\end{align}\]

ganeshie8 (ganeshie8):

Note that in the given interval all trig functions are positive so you don't need to wry about canceling radicals

ganeshie8 (ganeshie8):

evaluating sec^3 should be pretty straightforward ?

OpenStudy (anonymous):

Yes. Thank you

OpenStudy (anonymous):

What would be the correct etiquette here? Who should get the medal from me? zepdrix lead me through a larger portion of the work, but ganeshie8 brought the problem closer to the answer

zepdrix (zepdrix):

We're not medal hogs, it shouldn't matter either way c: Other people gave some medals for this question also, so that's appreciated ^^

ganeshie8 (ganeshie8):

^^

OpenStudy (anonymous):

sorry, still can't get this one...